Shinsuke Uno walked his dog on a path through a park which formed a right triangle. One leg of the right triangle was 7 meters longer than the other leg. The hypotenuse of the right triangle was 13 meters long. How far did he walk altogether?
Answer by asuar010(338)

First set up the Pythagorean theorem as x^2 + (x+7)^2 =13^2
then start solving: x^2+x^2+14x+49=169 then 2x^2+14x-120=0 then factor it as
(2x-10)*(x+12)=0 then 2x-10=0 and x+12=0 solve for x in each case:
x=-12 and x=5, and since a distance cannot be negative the only possible answer
is 5 so the total distance traveled is 5+12+13=30